Main INDEX
Monthly INDEX
PREV
NEXT
User name jpchen
Log entry time 15:08:41 on September27,2006
Entry number 185591
keyword=day shift summary
Day shift summary
shift workers: J. P. Chen, E. Khrosinkova, O. Hansen.
8:00 in the middle of changing to kin1
Left angle changed to 24 degree.
Both spectrometer momentum to 0.66369 GeV/c, left settled.
Right NMR not locking, cummunication problem. Heidi came in, tried reboot, not working.
9:05 requested controlled access.
9:20 ARM arrived, Heidi, Gary, Xiaodong, Ron Guy went to the hall.
9:40 Per Ron Guy's request, closed 1.5" carbon door.
9:45 Heidi and Gary came out. Right NMR works now.
9:55 Xiaodong and Ron Guy came out the hall. Requested beam.
10:08 Beam arrived, started production
11:22 HRSR to 17 degrees. Raster on. Asking for 2 uA.
11:25 Asking for beam position to be zeroed at target, is off by 0.5mm.
11:35 Resuming production, C target, 2 uA, raster on, kinematics 1.
11:55 Moving to Ta target, asking for 2 uA.
12:15 Moving to 4cm LD2 target.
13:35 Moving to 4cm LH2 target.
14:20 Moving to 4cm dummy target.
14:50 Completed data taking for ED.
requested controlled access for configuration change,
which will last to the end of the shift.
Run 5247/25247 10:08, 2 uA on C, 1M/0.6M, Kin1, P=-0.66369, t_L=t_R=24 degree, Raster on
Run 5248/25248 10:36, 2 uA on C, 1M/0.6M, kin1, raster off.
run 5249/25249 11:35, 2 uA on C, 1M/0.8M, kin2, t_R=17 deg, raster on.
run 5250/25250 12:00, 2uA on Ta, 0.13M/0.5M, kin2.
run 5251/25251 12:28, 1uA on LD2, 1M/~0.9M, kin2.
run 5252/25252 12:45, 1uA on LD2, 1M/0.95M, kin2.
run 5253/25253 13:18, 1uA on LD2, 1M/0.97M, kin2.
run 5254/25254 13:40, 1uA on LH2, 1M/0.91M, kin2.
run 5255/25255 13:57, 1uA on LH2, 1M/1M, kin2.
run 5256/25256 14:13, 1uA on LH2, 0.3M/0.3M, kin2.
run 5257/25257 14:31, 1uA on 4 cm dummy, 1M/1.1M, kin2.